# Simplex volumes in hyperplane arrangements

### v1

@@ 3 Minimum volume $d$-simplices
Kobon Fujimura asked in his book (called ”The Tokyo puzzle” [18]) that what is the largest number $K(n)$ of non-overlapping triangles whose sides lie on an arrangement of $n$ lines (Grünbaum [24] had studied a similar problem slightly earlier than Fujimura, but only in the case of arrangements in the projective plane). A first upper bound was established by Saburo Tamura, which is $K(n)\leq\lfloor\frac{n(n-2)}{3}\rfloor$. In 2007, Clément and Bader [6] found a slightly improved upper bound:
@@
For even $n$, Bartholdi et al. [1] provided a slightly sharper upper bound of $\lfloor\frac{n}{3}(n-\frac{7}{3})\rfloor$. On the other hand, Füredi and Palásti [19] having constructed an arrangement of $n$ lines which contains $\lfloor\frac{n(n-3)}{3}\rfloor$ non-overlapping triangles, which gives $K(n)\geq\lfloor\frac{n(n-3)}{3}\rfloor$.
We now consider a higher-dimensional analogue. Let $K_{d}(n)$ be the largest number of nonoverlapping $d$-dimensional simplices whose facets lie on an arrangement of $n$ hyperplanes in $\mathbb{R}^{d}$. Note that $K(n)=K_{2}(n)$.
@@
The proof is presented in the appendix. For $d=2$, it immediately follows that $m_{2}(n)\leq K_{2}(n)$ from the fact that cutting a triangle with a single line always creates a smaller triangle in its interior ([9]). However, unfortunately, for $d\geq 3$, the same argument does not hold. Indeed, a $d$-simplex cut by a hyperplane creates a new $d$-simplex with volume smaller than the original in the interior only if the hyperplane separates one vertex from all the others. See Figure 1 for the case $d=3$. We provide an upper bound for $m_{d}(n)$ by a different approach.
Figure 1: Left: A cut that creates a $3$-simplex in the interior of a 33-simplex. Right: A cut that does not create a $3$-simplex in the interior.
@@
Upper bound: We begin by demonstrating the three-dimensional case. Let $\mathcal{H}$ be an arrangement of $n$ planes in $\mathbb{R}^{3}$, and let $V_{0}$ be the volume of a tetrahedron with minimum volume formed by $\mathcal{H}$. We denote by $\tau(H_{1},H_{2},H_{3},H_{4})$ the tetrahedron determined by $H_{1},H_{2},H_{3},H_{4}$, and let $\tilde{H_{i}}\subset H_{i}\,\,(i=1,2,3,4)$ be its faces.
Consider $H_{1},H_{2},H_{3}\in\mathcal{H}$ intersecting at a single point. These divide $\mathbb{R}^{3}$ into $8$ connected components. Let $R$ be one of them. Then by Proposition 2.1, there exists a connected branch $\sigma_{R}$ of a surface in the region $R$ such that for any point $p$ on $\sigma_{R}$, the tangent plane $T_{p}(\sigma_{R})$ satisfies that the tetrahedron $\tau(H_{1},H_{2},H_{3},T_{p}(\sigma_{R}))$ has volume $V_{0}$. Moreover, by fixing a coordinate system $(x,y,z)$ and applying an appropriate affine transformation, we may assume $H_{1}=\{z=0\}$, $H_{2}=\{y=0\}$, $H_{3}=\{x=0\}$. Then the connected branch corresponding to $\sigma_{R}$ is $C_{R}:xyz=2V_{0}/9\,\,(x,y,z>0)$. For a point $p$ on $C_{R}$, let $T_{p}(C_{R})$ denote its tangent plane.
Claim 1.
$|\{p\in C_{R}:T_{p}(C_{R})\in\mathcal{H}\}|\leq 1$, that is, there is at most one $H\in\mathcal{H}$ such that the tetrahedron $\tau(H_{1},H_{2},H_{3},H)$ has a minimum volume in $R$.
Proof.
[…]
Figure 2:
Suppose there are two planes $T,T^{\prime}\in\mathcal{H}$ such that for some $p,p^{\prime}\in C_{R}$, we have $T=T_{p}(C_{R}),\,\,T^{\prime}=T_{p^{\prime}}(C_{R})$. Then the interior of the face $\tilde{T}$ of the tetrahedron $\tau=\tau(H_{1},H_{2},H_{3},T)$ intersects $T^{\prime}$. Without loss of generality, we may assume that $T^{\prime}$ does not intersect only $H_{3}\cap\tilde{T}$ among the three edges of $\tilde{T}$. Let $s_{i}$ denote the intersection point of $T^{\prime}$ with $H_{i}\cap T$ for $i\in\{1,2\}$. The vertex $H_{1}\cap H_{2}\cap T$ of $\tau$ can be represented as $S=(s,0,0)$ for some $s>0$. Also, the vertex $H_{1}\cap H_{2}\cap H_{3}$ of $\tau$ is the origin $O$. From the positions of $s_{1}$ and $s_{2}$, it follows that $T^{\prime}$ must intersect the open half-line $\{(t,0,0):t>0\}(\subset H_{1}\cap H_{2})$. Let $H_{1}\cap H_{2}\cap T^{\prime}=\{(u,0,0)\}$. We may assume $u>s$. Indeed, otherwise the tetrahedron $Sus_{1}s_{2}$ would be contained in the interior of $\tau$, contradicting the minimality of the volume of $\tau$. When $u>s$, $T^{\prime}$ intersects $\tilde{H_{1}}\cap\tilde{H_{3}}$ at some point $t_{1}$ and $\tilde{H_{2}}\cap\tilde{H_{3}}$ at some point $t_{2}$. Then the tetrahedron $Out_{1}t_{2}$ contains the tetrahedron $Sus_{1}s_{2}$ in its interior. However, the tetrahedron $Out_{1}t_{2}$ is $\tau(H_{1},H_{2},H_{3},T^{\prime})$, contradicting the minimality of volume. See Figure 2. ∎
By the same argument for the other connected components besides $R$, we find that there are at most $8$ planes $H\in\mathcal{H}$ such that $\tau(H_{1},H_{2},H_{3},H)$ has minimum volume $V_{0}$. Therefore, the number of tetrahedra with minimum volume in $\mathcal{H}$ is at most $8\binom{n}{3}$.
It turns out that an even simpler approach works for higher dimensions as follows. As in the three-dimensional case, we consider hyperplanes $H_{1}=\{x_{1}=0\},\ldots,H_{d}=\{x_{d}=0\}$ in $\mathbb{R}^{d}$ and a connected branch $C_{R}$ of a hypersurface of degree $d$ belonging to the region $R=\mathbb{R}_{\geq 0}^{d}$ which satisfy the property of Proposition 2.1 with minimum volume. We fix two tangent hyperplanes $T,T^{\prime}$ of $C_{R}$, and denote the $d$-simplices they form by $\tau=\tau(H_{1},\ldots,H_{d},T)$ and $\tau^{\prime}=\tau(H_{1},\ldots,H_{d},T^{\prime})$, respectively. Now, by a well-known result of Shannon [35], the arrangement of $d+2$ hyperplanes $\{H_{1},\ldots,H_{d},T,T^{\prime}\}$ has two $d$-simplicial cells $\sigma_{1},\sigma_{2}$. Clearly, $T^{\prime}$ must passes through the interior of $\tau$ and divides $\tau$ into two regions $R_{1},R_{2}$. We may assume that neither $R_{1}$ nor $R_{2}$ is a $d$-simplicial cell (otherwise this would contradict the minimality of the volume of $\tau$). Therefore, neither $\sigma_{1}$ nor $\sigma_{2}$ intersects $\tau$. Then, $\tau^{\prime}$ can be expressed as the union of one of $R_{1},R_{2}$ and one of $\sigma_{1},\sigma_{2}$. This contradicts the minimality of the volume of $\tau^{\prime}$. Hence, the number of minimum $d$-volume $d$-simplices is at most $2^{d}\binom{n}{d}$.
@@
Lower bound: Let us denote by $S_{d}$ the set of all permutations of $\{1,\ldots,d\}$ and let $\pi\in S_{d}$ be a permutation. The Coxeter–Freudenthal–Kuhn (CFK) triangulation of the unit hypercube $0,1]^{d}$ is a triangulation of $0,1]^{d}$ into $d$-simplices $\sigma_{\pi}=[v_{0}(\pi),\ldots,v_{d}(\pi)]$, whose vertices are given by
@@
where $e_{1},\ldots,e_{d}$ is a standard basis of $\mathbb{R}^{d}$. Note that the points $v_{0}(\pi),\ldots,v_{d}(\pi)\in\mathbb{R}^{d}$ are vertices of $0,1]^{d}$ and the CFK triangulation decomposes $0,1]^{d}$ into $d!$ $d$-simplices with volume $1/d!$. It is easy to observe that this decomposition is realized by $\binom{d}{2}$ hyperplanes
$\{x_{p}-x_{q}=0:1\leq p<q\leq d\}.$
See Figure 3 and refer to [8], [17], [28] for further details
@@
Now, we consider $d(n+1)$ hyperplanes given by
$x_{i}=k\quad(i\in\{1,\ldots,d\},\,\,k\in\{0,\ldots,n\}),$
which form an arrangement that decomposes the hypercube $0,n]^{d}$ into $n^{d}$ unit hypercubes. Furthermore, by adding $\binom{d}{2}(2n-1)$ hyperplanes
@@
We can see that each unit hypercube admits the CFK triangulation (this is called K1 CFK triangulation). The size of the resulting hyperplane arrangement is
@@
and the number of (minimum volume) $d$-simplices is
$d!\cdot n^{d},$
which gives us a lower bound $\Omega_{d}(n^{d})$.
@@ 4 Maximum volume tetrahedra
In this section, we show a lower bound for $M_{3}(n)$. A linear lower bound for $M_{3}(n)$ can be easily obtained using $3$ planes together with n3n-3 tangent planes to a hypersurface having them as asymptotic planes. By generalizing the argument in [9] to 3 dimensions, we obtain the following lower bound. See 1.4 The proposition they proved for their construction giving $M_{2}(n)>6n/5-O(1)$ can be extended as follows.
@@
If there are no parallel planes in the arrangement, then we can find a rectangular box $ABCDEFGH$ such that adding any plane to the arrangement that intersects the segments $AB$, CDCD, EFEF and $GH$ creates no new maximum volume tetrahedra.
@@
Proposition 4.2 (Generalization of Proposition 4.2 in [9]).
If $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are arrangements of planes which contains no parallel planes, then there exists affine transformations $\varphi,\psi$ such that
$T(\varphi(\mathcal{H}_{1})\cup\psi(\mathcal{H}_{2}))\geq T(\mathcal{H}_{1})\cup T(\mathcal{H}_{2})+2.$
@@
In an arrangement $\mathcal{H}_{1}$, there exists a rectangular box satisfying Proposition 4.1 with vertices $A_{1},B_{1},C_{1},D_{1},E_{1},F_{1},G_{1},H_{1}$, and similarly in $\mathcal{H}_{2}$, there exists one with vertices $A_{2},B_{2},C_{2},D_{2},E_{2},F_{2},G_{2},H_{2}$. By applying appropriate affine transformations to each, we can arrange $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ so that the volumes of their respective maximum-volume tetrahedra are equal and the two rectangular boxes cross each other as shown in the Figure 5. Let $\varphi$ denote the affine transformation applied to $\mathcal{H}_{1}$ and $\psi$ the affine transformation applied to $\mathcal{H}_{2}$.
@@
We observe that the plane arrangement $\mathcal{S}$ forming a star-shaped badge as shown in the Figure 5 contains $6$ planes and $5$ tetrahedra of maximum volume. Starting from $\mathcal{S}$, we repeatedly add copies of $\mathcal{S}$ so as to satisfy the above property. This allows us to construct a planar arrangement of $6k$ planes containing 7k27k-2 tetrahedra of maximum volume for $k\geq 1$.
We are not aware of any construction providing a non-trivial lower bound for $M_{d}(n)$ for $d\geq 4$.
@@ 5 Hyperplanes defining distinct volume $d$-simplices
Let $n]=\{1,\ldots,n\}$ and denote $r_{k}(n)$ to be the size of the largest $S\subseteq[n]$ such that $S$ contains no $k$ elements in arithmetic progression ($k$-AP). Here is the key proposition.
@@
Determining the value of $r_{k}(n)$ is one of the fundamental problems in arithmetic combinatorics, and both upper and lower bounds have been progressively improved by many researchers. In 2009, Green and Tao [22, 23] proved that there exists a constant $c>0$ such that $r_{4}(n)\ll n(\log n)^{-c}$. For $k\geq 4$, Leng et al. [29] recently obtained an upper bound $r_{k}(n)<<ne^{-(\log\log n)^{c_{k}}}$ for some $c_{k}\in(0,1)$. See [41, 20, 21] for further details. Theorem 1.5 can be immediately derived from these bounds and Proposition 5.1. Therefore, any improvement of upper bound for $r_{k}(n)$ automatically improves that for $D_{d}(n)$.
@@
We generalize this construction to higher dimensions. Assume $d$ is even. Write $\mathbb{R}^{d}$ as the orthogonal sum of a two-dimensional subspace, thus
@@
Now let $i_{1}<i_{2}<\cdots<i_{d+2}$ be a $(d+2)$-arithmetic progression in $[n]$ with common difference $L$. As the arrangement $\mathcal{H}$ is in general position, the hyperplanes $\{H_{i_{1}},\ldots,H_{i_{d+1}}\}$ form a unique $d$-simplex $S$, and the hyperplanes $\{H_{i_{2}},\ldots,H_{i_{d+2}}\}$ form a unique $d$-simplex $S^{\prime}$. Since $T_{L}(H_{i_{k}})=H_{i_{k+1}}$, we have $T_{L}(S)=S^{\prime}$ which means that $S$ and $S^{\prime}$ have the same volume. Therefore, by the definition of $r_{k}(n)$, every subfamily of $\mathcal{H}$ of size (at least) $r_{d+2}(n)+1$ always contain two $d$-simplices with the same volume.
Now, we assume $d$ is odd, we decompose $\mathbb{R}^{d}$ into an orthogonal direct sum of (d1)/2(d-1)/2 two-dimensional vector spaces and one one-dimensional vector space as follows:
@@
Observe that $T_{D}(P_{i,j})=P_{i+D,j},\,T_{D}((0,\ldots,0,i))=(0,\ldots,0,i+D)$ (for $i+D\leq n$), hence $T_{D}(H_{i})=H_{i+D}.$ The same argument as in the previous case shows that any subfamily of $\mathcal{H}$ with size $r_{d+2}(n)+1$ contains two $d$-simplices with the same volume.
@@
As previouly mentioned, Damásdi et al. [9] obtained the lower bound with some additional assumptions: if there are no six lines that are tangent to a common conic (observe that $5$ lines always have a common tangent conic), then an arrangement of $n$ lines always contains a subset of size $n^{1/5}$ whose induced triangles have all distinct volumes. Their argument can be applied to general dimensions, and the following can be shown. If there are no $d(d+3)/2$ hyperplanes that are tangent to a common hypersurface, then then an arrangement of $n$ hyperplanes always contains a subset of size $n^{\frac{1}{2d+1}}$ whose induced $d$-simplices have all distinct volumes. On the other hand, We are not aware of any lower bounds for $D_{d}(n)$ for $d\geq 2$.
@@
holds for all $d\geq 3$. Suppose there are two parallel hyperplanes in an arrangement. If we perturb one of them by slightly rotating its normal vector, the number of $d$-simplicial cells in the arrangement does not decrease. Hence, we may assume that no two hyperplanes are parallel.
Consider an arrangement $\mathcal{H}$ of $n$ hyperplanes, and fix one hyperplane $H$ of $\mathcal{H}$. Then, restricting the remaining to $H$ yields an arrangement $\mathcal{S}$ of (d2)(d-2)-dimensional hyperplane in $H(\cong\mathbb{R}^{d-1})$. Since no two in $\mathcal{H}$ are parallel, no two in $\mathcal{S}$ are parallel either. The maximum number of (d1)(d-1)-simplicial cells in $\mathcal{L}$ is $K_{d-1}(n-1)$. Therefore, the total number of (d1)(d-1)-simplicial cells in $\mathcal{A}$ is at most $nK_{d-1}(n-1)$.
We now divide into two cases. First assume that no $d+1$ hyperplanes intersect at a single point. In this case, No two $d$-simplicial cells share a common (d1)(d-1)-simplicial face. Otherwise, at least one of the $d+1$ vertices of the common face intersects at least $d+1$ hyperplanes. See Figure 7. Therefore,
@@
Next, assume that $d+1$ (or more) hyperplanes intersect at a single point. If a $(d+1)$-th hyperplane is added through an existing intersection point, the number of (d1)(d-1)-simplicial faces decreases by $d+1$. On the other hand, two $d$-simplicial cells sharing a common (d1)(d-1)-simplicial face can be created only by adding a new hyperplane passing through an existing point. See Figure 7. Therefore,
@@
Figure 6:
Figure 7: In the case $d=3$: If four planes intersect at a single point, adding a fourth plane through an existing point decreases the number of triangles by four, with at most two of them are saved.

### v2

@@ 3 Minimum volume $d$-simplices
Kobon Fujimura asked in his book (called ”The Tokyo puzzle” [Fujimura1978]) what is the largest number $K(n)$ of non-overlapping triangles whose sides lie on an arrangement of $n$ lines (Grünbaum [grunbaum1972arrangements] had studied a similar problem slightly earlier than Fujimura, but only in the case of arrangements in the projective plane). A first upper bound was established by Saburo Tamura, which is $K(n)\leq\lfloor\frac{n(n-2)}{3}\rfloor$. In 2007, Clément and Bader [clement2007tighter] found a slightly improved upper bound:
@@
For even $n$, Bartholdi et al. [bartholdi2008simple] provided a slightly sharper upper bound of $\lfloor\frac{n}{3}(n-\frac{7}{3})\rfloor$. On the other hand, Füredi and Palásti [furedi1984arrangements] having constructed an arrangement of $n$ lines which contains $\lfloor\frac{n(n-3)}{3}\rfloor$ non-overlapping triangles, which gives $K(n)\geq\lfloor\frac{n(n-3)}{3}\rfloor$.
We now consider a higher-dimensional analogue. Let $K_{d}(n)$ be the largest number of nonoverlapping $d$-simplices whose facets lie on an arrangement of $n$ hyperplanes in $\mathbb{R}^{d}$. Note that $K(n)=K_{2}(n)$.
For $d=2$, it immediately follows that $m_{2}(n)\leq K_{2}(n)$ from the fact that cutting a triangle with a single line always forms a smaller triangle in its interior. However, unfortunately, the same argument does not hold for $d\geq 3$. Indeed, a $d$-simplex cut by a hyperplane creates a new $d$-simplex with $d$-volume smaller than the original in the interior if and only if the hyperplane separates one vertex from all the others. See Figure 1 for the case $d=3$.
Figure 1: Left: a cut creating a $3$-simplex in the interior of a 33-simplex. Right: a cut not creating a $3$-simplex in the interior.
Nevertheless, finding bounds for $K_{d}(n)$ might be an interesting problem for its own sake.
@@
The proof is presented in the Appendix. We provide an upper bound for $m_{d}(n)$ by a different approach.
@@
Upper bound: Let $\mathcal{H}=\{H_{1},\dots,H_{n}\}$ be an arrangement of hyperplanes in $\mathbb{R}^{d}$, and let $m_{0}$ be the minimum $d$-volume of a $d$-simplex determined by $\mathcal{H}$. Assume that there exists a $d$-tuple $(H_{i_{1}},\dots,H_{i_{d}})$ of hyperplanes in $\mathcal{H}$ meeting at exactly one point. Applying Proposition 2.1 to $(H_{i_{1}},\dots,H_{i_{d}})$ with $V_{0}=m_{0}$, any hyperplane $T$ such that $\{H_{i_{1}},\dots,H_{i_{d}},T\}$ forms a $d$-simplex of $d$-volume $m_{0}$ must be tangent to one of the two fixed algebraic hypersurfaces of degree $d$. These two hypersurfaces have $2^{d}$ connected components in total, and each of the $2^{d}$ connected components determined by $H_{i_{1}},\dots,H_{i_{d}}$ contains exactly one of them.
Let $C_{R}$ be the unique connected component contained in a fixed connected component $R$ in $\{H_{i_{1}},\dots,H_{i_{d}}\}$ . It suffices to show that at most one hyperplane in $\mathcal{H}\setminus\{H_{i_{1}},\dots,H_{i_{d}}\}$ is tangent to $C_{R}$. For a contradiction, we assume that there exist two distinct hyperplanes $T,T^{\prime}\subset\mathcal{H}\setminus\{H_{i_{1}},\dots,H_{i_{d}}\}$ tangent to $C_{R}$. Let $\tau$ be the $d$-simplex formed by $H_{i_{1}},\dots,H_{i_{d}},T$, and let $\tau^{\prime}$ be the $d$-simplex formed by $H_{i_{1}},\dots,H_{i_{d}},T^{\prime}$. Note that $\tau$ and $\tau^{\prime}$ are $d$-simplices of $d$-volume $m_{0}$.
Now consider the arrangement $\mathcal{A}=\{H_{i_{1}},\dots,H_{i_{d}},T,T^{\prime}\}$. A well-known result of Shannon [shannon1979simplicial] states that an arrangement of $N(\geq d+1)$ hyperplanes in general position in $\mathbb{R}^{d}$ has NdN-d simplicial $d$-cells, and hence $\mathcal{A}$ has two simplicial $d$-cells. Observe that these two simplicial $d$-cells have $T$ as a facet, and they do not lie on the same side of $T$. Therefore, at least one of them is strictly contained in $\tau$ or $\tau^{\prime}$. This contradicts the minimality of $m_{0}$. Hence, we have $m_{d}(n)\leq 2^{d}\binom{n}{d}=O_{d}(n^{d})$.
@@
Lower bound: Denote by $S_{d}$ the set of all permutations of $\{1,\ldots,d\}$ and let $\pi\in S_{d}$ be a permutation. The Coxeter–Freudenthal–Kuhn (CFK) triangulation of the unit hypercube $0,1]^{d}$ is a triangulation of $0,1]^{d}$ into $d$-simplices $\sigma_{\pi}=[v_{0}(\pi),\ldots,v_{d}(\pi)]$, whose vertices are given by
@@
where $e_{1},\ldots,e_{d}$ is a standard basis of $\mathbb{R}^{d}$. Note that the points $v_{0}(\pi),\ldots,v_{d}(\pi)\in\mathbb{R}^{d}$ are vertices of $0,1]^{d}$ and the CFK triangulation decomposes $0,1]^{d}$ into $d!$ $d$-simplices with $d$-volume $1/d!$. It is easy to observe that this decomposition is realized by $\binom{d}{2}$ hyperplanes $\{x_{p}-x_{q}=0:1\leq p<q\leq d\}.$ See Figure 2 and refer to [coxeter1934discrete], [freudenthal1942simplizialzerlegungen], [kuhn1968simplicial] for further details
@@
Now, we consider $d(n+1)$ hyperplanes given by $x_{i}=k\quad(i\in\{1,\ldots,d\},\,\,k\in\{0,\ldots,n\}),$ which form an arrangement that decomposes the hypercube $0,n]^{d}$ into $n^{d}$ unit hypercubes. Furthermore, by adding $\binom{d}{2}\cdot(2n-1)$ hyperplanes
@@
We can see that each unit hypercube admits the CFK triangulation (this is called K1 triangulation [REIA2025104237]). The size of the resulting hyperplane arrangement is
@@
and the number of (minimum volume) $d$-simplices is $d!\cdot n^{d},$ which gives us a lower bound $\Omega_{d}(n^{d})$.
@@ 4 Maximum volume tetrahedra
In this section, we show a lower bound for $M_{d}(n)$. A linear lower bound $M_{d}(n)\geq n-d$ can be easily obtained as follows: Fix coordinates $(x_{1},\dots,x_{d})$ in $\mathbb{R}^{d}$ and let $H_{i}:=\{x_{i}=0\}$, $i=1,\dots,d$ and consider the branch
$\sigma:=\{x\in\mathbb{R}^{d}:\ \prod_{i=1}^{d}x_{i}=1,\ x_{1},\ldots,x_{d}>0\}.$
Choose ndn-d distinct hyperplanes $T_{1},\dots,T_{n-d}$ tangent to $\sigma$, with their tangent points contained in a sufficiently small neighborhood of some fixed point $u\in\sigma$. For each $j$, the hyperplanes $H_{1},\dots,H_{d},T_{j}$ bound a $d$-simplex in the orthant $\{(x_{1},\ldots,x_{d}):x_{1},\ldots,x_{d}\geq 0\}$, and its $d$-volume equals $d^{d}/d!$. Since the points of tangency of $T_{1},\dots,T_{n-d}$ are chosen sufficiently close to each other, any bounded $d$-simplex of the arrangement that uses at least two of the hyperplanes $T_{j}$ has $d$-volume smaller than $d^{d}/d!$. Hence the simplices bounded by $H_{1},\dots,H_{d}$ and one $T_{j}$ are of maximum $d$-volume in the arrangement, and therefore $M_{d}(n)\geq n-d$.
4.1 Explicit lower bound construction for $M_{3}(n)$
As mentioned in the introduction, Damasdi et al. [damasdi2020triangle] obtained the lower bound $M_{2}(n)>\frac{6}{5}n-O(1)$. The idea of their lower bound construction is the following: They begin with a family of five lines forming a regular pentagon, and then recursively place arrangements together. At each step, two arrangements are combined by applying a suitable affine transformation so that the areas of their maximum area triangles coincide, and placing them in such a way that the maximum area does not increase. Moreover, by sliding one of the whole families of lines, one can create two additional maximum area triangles. Consequently, we obtain a construction in which the number of maximum area triangles increases by seven whenever a family of five lines is added. We see that this idea can also work in three dimensions.
@@ 4.1 Explicit lower bound construction for $M_{3}(n)$
$M_{3}(n)>\frac{7}{6}n-O(1).$
The following can be proved using the same procedure as in their proof. Details are left to the reader.
Proposition 4.2.
If there are no two parallel planes in the arrangement, then we can find a rectangular box $ABCDEFGH$ such that adding any plane to the arrangement that intersects the segments $AB$, CDCD, EFEF and $GH$ creates no new maximum volume tetrahedra.
@@
Proposition 4.3 ($3$-dimensional analogue of Proposition 4.2 in [damasdi2020triangle]).
If $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are arrangements of planes neither of which contains two parallel planes, then there exist affine transformations $\varphi,\psi$ such that
$T(\varphi(\mathcal{H}_{1})\cup\psi(\mathcal{H}_{2}))\geq T(\mathcal{H}_{1})+T(\mathcal{H}_{2})+2.$
@@
In an arrangement $\mathcal{H}_{1}$, there exists a rectangular box satisfying Proposition 4.2 with vertices $A_{1},B_{1},C_{1},D_{1},E_{1},F_{1},G_{1},H_{1}$, and similarly in $\mathcal{H}_{2}$, there exists one with vertices $A_{2},B_{2},C_{2},D_{2},E_{2},F_{2},G_{2},H_{2}$. By applying appropriate affine transformations to each, we can arrange $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ so that the volumes of their respective maximum volume tetrahedra are equal and the two rectangular boxes cross each other as shown in Figure 4. Let $\varphi$ denote the affine transformation applied to $\mathcal{H}_{1}$ and $\psi$ the affine transformation applied to $\mathcal{H}_{2}$.
@@
We observe that the arrangement of planes $\mathcal{S}\subset\mathbb{R}^{3}$ forming a star-shaped badge as shown in Figure 4 consists of six planes and forming five maximum volume tetrahedra. Starting from $\mathcal{S}$, we repeatedly add suitably transformed copies of $\mathcal{S}$ so that the above proposition applies at each step. Consequently, for every $k\geq 1$ we obtain an arrangement of $6k$ planes in $\mathbb{R}^{3}$ that determines at least 7k27k-2 maximum volume tetrahedra, which proves Theorem 1.4.
4.2 General lower bound for $M_{d}(n)$
We improve here the lower bound achieved by Proposition 4.1 and also provides lower bounds for all $d\geq 3$.
See 1.4
The Theorem follows immediately from the following observation.
Observation 1.
For all $d\geq 3$, we have
$M_{d}(n)\geq M_{d-1}(n-1).$
Proof.
[…]
We are not aware of any upper bound for $M_{d}(n)$ that holds for all $d\geq 3$. Moreover, we do not have any reason to believe that the correct upper bound is also close to linear. Whether $M_{d}(n)$ is linear in $n$ is open even in the case $d=2$. Note that Damásdi et al. [damasdi2020triangle] conjectured that $M_{2}(n)=O(n^{1+\varepsilon})$ for every $\varepsilon>0$.
@@ 5 Hyperplanes defining distinct volume $d$-simplices
Let $n]=\{1,\ldots,n\}$ and denote by $r_{k}(n)$ the size of the largest $S\subseteq[n]$ such that $S$ contains no $k$ elements in arithmetic progression. The following is the key proposition.
@@
Determining the value of $r_{k}(n)$ is one of the fundamental problems in arithmetic combinatorics, and both upper and lower bounds have been progressively improved. In 2009, Green and Tao [green2006new, green2017new] proved that there exists a constant $c>0$ such that $r_{4}(n)\ll n(\log n)^{-c}$. For $k\geq 4$, Leng et al. [leng2024improved] recently obtained an upper bound $r_{k}(n)<<ne^{-(\log\log n)^{c_{k}}}$ for some $c_{k}\in(0,1)$. See [szemeredi1975sets, gowers1998new, gowers2000arithmetic] for further details. Theorem 1.5 can be immediately derived from these bounds and Proposition 5.1. Therefore, any improvement of upper bound for $r_{k}(n)$ automatically improves that for $D_{d}(n)$.
@@
We generalize this construction to higher dimensions. First, assume that $d$ is even. Write $\mathbb{R}^{d}$ as the orthogonal sum of two-dimensional subspaces:
@@
Now let $i_{1}<i_{2}<\cdots<i_{d+2}$ be a $(d+2)$-arithmetic progression in $[n]$ with common difference $L$. As the arrangement $\mathcal{H}$ is in general position, the hyperplanes $\{H_{i_{1}},\ldots,H_{i_{d+1}}\}$ form a unique $d$-simplex $S$, and the hyperplanes $\{H_{i_{2}},\ldots,H_{i_{d+2}}\}$ form a unique $d$-simplex $S^{\prime}$. Since $T_{L}(H_{i_{k}})=H_{i_{k+1}}$, we have $T_{L}(S)=S^{\prime}$ which means that $S$ and $S^{\prime}$ have the same $d$-volume. Therefore, by the definition of $r_{k}(n)$, every subfamily of $\mathcal{H}$ of size (at least) $r_{d+2}(n)+1$ always contain two $d$-simplices with the same $d$-volume.
Now, we assume $d$ is odd, we decompose $\mathbb{R}^{d}$ into an orthogonal direct sum of (d1)/2(d-1)/2 2-dimensional vector spaces and one 1-dimensional vector space:
@@
Observe that $T_{D}(P_{i,j})=P_{i+D,j},\,T_{D}((0,\ldots,0,i))=(0,\ldots,0,i+D)$ (for $i+D\leq n$), hence $T_{D}(H_{i})=H_{i+D}.$ The same argument as in the previous case shows that any subfamily of $\mathcal{H}$ with size $r_{d+2}(n)+1$ contains two $d$-simplices with the same $d$-volume.
@@
As previouly mentioned, Damásdi et al. [damasdi2020triangle] obtained the lower bound with some additional assumptions: if there are no six lines that are tangent to a common conic (observe that five lines always have a common tangent conic), then an arrangement of $n$ lines always contains a subset of size $\Omega(n^{1/5})$ whose induced triangles have all distinct $d$-volumes. Their argument can be applied to general dimensions, and the following can be shown: If there are no $\binom{2d}{d}$ hyperplanes that are tangent to a common hypersurface, then then an arrangement of $n$ hyperplanes always contains a subset of size $\Omega_{d}(n^{\frac{1}{2d+1}})$ whose induced $d$-simplices have all distinct volumes. However, We are not aware of any lower bounds for $D_{d}(n)$ for $d\geq 2$.
@@
holds for all $d\geq 3$. Suppose there are two parallel hyperplanes in an arrangement. If we perturb one of them by slightly rotating its normal vector, the number of simplicial $d$-cells in the arrangement does not decrease. Hence, we may assume that no two hyperplanes are parallel.
Consider an arrangement $\mathcal{H}$ of $n$ hyperplanes, and fix one hyperplane $H$ of $\mathcal{H}$. Then, restricting the remaining to $H$ yields an arrangement $\mathcal{S}$ of n1n-1 (d2)(d-2)-flats in $H(\cong\mathbb{R}^{d-1})$. Since no two of hyperplanes in $\mathcal{H}$ are parallel, no two in $\mathcal{S}$ are parallel either. Since the number of simplicial (d1)(d-1)-cells in $\mathcal{S}$ is at most $K_{d-1}(n-1)$, the total number of simplicial (d1)(d-1)-cells in $\mathcal{H}$ is at most $nK_{d-1}(n-1)$.
We now divide into two cases. First assume that no $d+1$ hyperplanes intersect at a single point. Then, no two simplicial $d$-cells share a common (d1)(d-1)-face. Otherwise, at least one of the $d$ vertices of the common face intersects at least $d+1$ hyperplanes. See Figure 6. Therefore,
@@
Next, assume that $d+1$ (or more) hyperplanes intersect at a single point. If a $(d+1)$-th hyperplane is added through an existing intersection point, the number of simplicial (d1)(d-1)-faces decreases by $d+1$. On the other hand, two simplicial $d$-cells sharing a common simplicial (d1)(d-1)-face can be created only by adding a new hyperplane passing through an existing point. See Figure 6. Therefore,
@@
Figure 5: The case $d=3$: if two tetrahedron share a common face, then at least one vertex of that face is contained in at least 44 planes.
Figure 6: The case $d=3$: if four planes intersect at a single point, adding a fourth plane through an existing point decreases the number of triangles by four, with at most two of them are saved.